OpenStudy (inkyvoyd):

Can anyone explain how to find the integral of sec(x) or csc(x) with respect to x?

OpenStudy (anonymous):

lalaly is the integral master today

OpenStudy (lalaly):

\[\int\limits{secxdx}=\int\limits{secx \times \frac{secx+tanx}{secx+tanx}}dx\]\[=\int\limits{\frac{\sec^2 x+secxtanx}{secx+tanx}dx}\] let u=secx+tanx so the integration becomes \[\int\limits{\frac{du}{u}}=\ln|u|+C\]\[=\ln|secx+tanx|+C\]

OpenStudy (lalaly):

same for csc(x) but u multiply by (cscx+cotx)/(cscx+cotx)

OpenStudy (lalaly):

lol satelite:)

OpenStudy (inkyvoyd):

Holy crap, no wonder I couldn't figure it out on my own. Are there any other cases where we do this?

OpenStudy (anonymous):

lalaly is correct

OpenStudy (anonymous):

here is another version. apparently this was a big deal in the 17th century http://en.wikipedia.org/wiki/Integral_of_the_secant_function